Let two cylindrical conductors connected in parallel, to which potential difference V=170V is applied. The two conductors are made of the same material as the first is 6 times the length of the second and 1/5 the radius of the second. The resistance of the second is 469 ohms. Determine the equivalent resistance


1
Expert's answer
2021-06-28T17:08:03-0400

Gives

V=170V

"l_1=6l_2"

"r_2=\frac{r_1}{5}"


We know that

"R_1=\frac{\rho l_1}{\pi d_1^2}"

"R_2=\frac{\rho l_2}{\pi d_2^2}"


"\frac{R_1}{R_2}=\frac{\frac{\rho l_1}{\pi d_1^2}}{\frac{\rho l_2}{\pi d_2^2}}"

Put value

"\frac{R_1}{R_2}=\frac{6}{25}"

"R_2=469\Omega"


"R_1=469\times \frac{6}{25}=112\Omega"


"R_{net}=R_1||R_2"

"R_{net}=\frac{R_1R_2}{R_1+R_2}"


"R_{net}=\frac{112\times469}{112+469}=90\Omega"